# Cubic* criticality emerging from quantum loop model on triangular lattice

### v2

@@ Abstract
Quantum loop and dimer models are archetypal examples of correlated systems with local constraints, whose generic solutions are difficult to obtain due to the lack of controlled methods to solve them in the thermodynamic limit. Yet, these solutions are of immediate relevance towards both statistical and quantum field theories, as well as the fast-growing experiments in Rydberg atom arrays and quantum moiré materials, where the interplay between correlation and local constraints gives rise to a plethora of novel phenomena. In a recent work Yan et al. (2022a), it was found via sweeping cluster quantum Monte Carlo (QMC) simulations and field theory analysis that the triangular lattice quantum loop model (QLM) hosts a rich ground state phase diagram with lattice nematic (LN), vison plaquette (VP) crystals, and the $\mathbb{Z}_{2}$ quantum spin liquid (QSL) close to the Rokhsar-Kivelson (RK) point. Here, we focus on the continuous quantum critical point separating the VP and QSL phases and demonstrate via both static and dynamic probes in QMC simulations that this transition is of the (2+1)d Cubic* universality, in which the fractionalized visons in QSL condense to give rise to the crystalline VP phase, while leaving their trace in the anomalously large anomalous dimension exponent and pronounced continua in the dimer and vison spectra compared with those at the conventional Cubic or O(3) quantum critical points.
@@
Introduction.— Recently, the ground state phase diagram of quantum loop model (QLM) on triangular lattice Moessner and Sondhi (2001a); Moessner et al. (2001a); Plat et al. (2015); Roychowdhury et al. (2015) is mapped out with the sweeping cluster quantum Monte Carlo (QMC) algorithm Yan et al. (2022a, 2019, 2021a); Yan (2022); Yan et al. (2021b, 2022b, 2023) (shown in Fig. 1). The physics revealed therein Yan et al. (2022a) are profound, such as the hidden vison plaquette (VP) crystal, which is invisible from dimer correlations and is sandwiched between the lattice nematic (LN) order and the $\mathbb{Z}_{2}$ quantum spin liquid (QSL) close to the RK point Rokhsar and Kivelson (1988); Moessner and Sondhi (2001b); Henley (2004). Another interesting aspect is the structure of the phase diagram when connected to finite temperature, which is expected to be richer compared to its square lattice loop or dimer model cousins Alet et al. (2005a); Dabholkar et al. (2022, 2023). But perhaps the most intriguing aspect related to the quantum criticality of the model is the (2+1)d Cubic* transition from the VP phase to the $\mathbb{Z}_{2}$ QSL. At this transition, the VP order parameter, which emerges from the underlying resonance of the dimer pairs, fractionalizes into the vison order parameter of the O(3)/Cubic CFT primary field. The condensation of these fractionalized excitations, in turn, leads to a strong enhancement of the scaling dimension of the VP order parameter at the transition in an unconventional manner Wang et al. (2017); Sun et al. (2018); Wang et al. (2018); Isakov et al. (2012); Wang et al. (2021, 2021).
@@
Therefore, our motivation in this work is to elucidate the precise nature of the intriguing and unconventional Cubic* quantum critical point (QCP) through both static and dynamic probes. We aim to achieve a comprehensive understanding of this QCP by combining field-theoretical interpretation with state-of-the-art QMC simulations. We find that this transition exhibits an anomalously large anomalous dimension when viewed through the correlation functions of the $t$-term (the resonance term in the QLM Hamiltonian, explained below). These correlation functions represent composite objects of the fractionalized vison and correspond to the rank-2 tensor (or tensorial magnetization) of the (2+1)d O(3)/Cubic universality Hasenbusch (2023a, b) with a large scaling dimension, approximately $\eta_{T}\approx 1.42$Ballesteros et al. (1996); Aharony (1973); Hasenbusch and Vicari (2011); Adzhemyan et al. (2019); Aharony et al. (2022); Pelissetto and Vicari (2002); Chester et al. (2021); Rong and Su (2023). On the other hand, if one measures the correlation of the vison operator, the observed anomalous dimension is consistent with the conventionally small values of $\eta\approx 0.04$ for (2+1)d O(3)/Cubic universality Ballesteros et al. (1996); Aharony (1973); Hasenbusch and Vicari (2011); Adzhemyan et al. (2019); Aharony et al. (2022); Pelissetto and Vicari (2002); Chester et al. (2021). This sharp contrast clearly reveals the unconventional nature of the Cubic* transition that separates the unconventional VP phase, which is hidden from dimer measurements, and the $\mathbb{Z}_{2}$ QSL, where visons are the anyonic particles of the underlying topological order.
@@
Figure 1: Quantum loop model on the triangular lattice and its phase diagram. (a) $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the triangular lattice primitive vectors. The $t$ and $V$ terms are the kinetic and potential terms in the Hamiltonian Eq. (1), respectively. (b) The transition between the LN and VP crystals is first order Yan et al. (2022a), while the transition between VP and the $\mathbb{Z}_{2}$ QSL is continuous and of (2+1)d Cubic* universality Yan et al. (2022a). The correlation functions of the VP and vison order parameters around the Cubic* QCP ($V_{c}=0.59(2)$) are shown in Figs. 2 and 3. The schematic plot in the VP phase is the real-space vison correlations with the red (grey) color conveys its positive (negative) value in each triangle. The darker color stands for larger abosulate values of the correlation functions. The schematic plot in the QSL phase shows two visons connected by a string (the green dashed line), which represents the path $P$ of the vison-vison correlation function $v_{\gamma}(0)v_{\gamma}(r)$$=$$(-1)^{N_{P}}$, with $N_{P}$ the number of dimers cut along $P$. Here we set the vison in the lower triangle $v_{1}(0)=1$ as the reference to fix the gauge.
@@
In addition to these purely theoretical motivations, the QLM that we studied here has been widely treated as the low-energy effective model for many frustrated magnets Moessner and Sondhi (2001a); Moessner et al. (2001a); Roychowdhury et al. (2015); Plat et al. (2015); Yan et al. (2021b, 2022b, 2023); Wang et al. (2017, 2017); Zhang et al. (2018); Feng et al. (2017); Wen (2017); Wei et al. (2020); Li et al. (2020); Hu et al. (2020); Wen et al. (2019); Feng et al. (2018); Wen and Lee (2019); Zhou et al. (2017); Broholm et al. (2020); Da Liao et al. (2021); Zhou et al. (2022); Moessner et al. (2001b); Ivanov (2004); Ralko et al. (2005); Misguich and Mila (2008); Ran et al. (2023) and blockaded cold-atom arrays  Samajdar et al. (2021); Verresen et al. (2021); Zhou et al. ; Samajdar et al. (2023); Semeghini et al. (2021) in condensed matter and cold atom experiments Semeghini et al. (2021); Scholl et al. (2021). In the Rydberg array, static characteristics can be obtained easily via snapshot technique Scholl et al. (2021); Semeghini et al. (2021); Tran et al. (2023); Wurtz et al. (2023), while dynamic information can be measured through the real-time evolution Surace et al. (2020); Notarnicola et al. (2020); Guardado-Sanchez et al. (2018); Turner et al. (2018). Similarly, static and dynamic information for quantum magnets can be detected by neutron scattering or nuclear magnetic resonance experiments Feng et al. (2017); Wei et al. (2020); Hu et al. (2020); Paddison et al. (2017); Zorko et al. (2017); Wen et al. (2019); Zeng et al. (2023), and our computational scheme of the QMC + stochastic analytic continuation (SAC) Sandvik (1998a); Beach (2004); Syljuåsen (2008); Shao and Sandvik (2023) for the frustrated spin model, QDM, and QLM models have provided consistent static and dynamic information that have been used to explain experiments Shao et al. (2017); Li et al. (2020); Hu et al. (2020); Sun et al. (2018); Ma et al. (2018); Wang et al. (2021, 2021); Zhou et al. (2021); Yan et al. (2021b, 2022b, 2023); Bera et al. (2022). Based on these previous experiences, in this work, our QMC static correlations reveal different scaling dimensions at the Cubic* QCP, corresponding to the different constituent operators in the CFT data for $\eta_{T}$ and $\eta$. At the same time, our QMC+SAC dynamic measurements exhibit continua of the dimer and vison spectra as the dynamic signature of the $\mathbb{Z}_{2}$ topological order and its associated vison condensation in the vicinity of the Cubic* QCP.
@@
where $\alpha$ denotes all the rhombi (with three orientations) on the triangular lattice as shown in Fig. 1(a). The local constraint of the fully packed QLM requires two dimers to touch every site in any configuration. The kinetic term is controlled by $t$, which generates the dimer pair resonance on every flippable plaquette while respecting the local constraint, and $V$ is the repulsion ($V>0$) or attraction ($V<0$) between dimers facing each other. The RK point is located at $V=t=1$ and has an exact $\mathbb{Z}_{2}$ QSL solution Moessner and Sondhi (2001a). We set $t=1$ as the energy unit and perform the simulations for system sizes $L=8,12,16,20,24$ with the inverse temperature $\beta=\frac{1}{T}=L$ via the sweeping cluster QMC methods Yan et al. (2019); Yan (2022); Yan et al. (2021b, 2022b, 2022), and utilize the SAC scheme Sandvik (1998b); Shao et al. (2017); Sun et al. (2018); Ma et al. (2018); Yan et al. (2021b); Shao and Sandvik (2023); Li et al. (2020); Hu et al. (2020); Wang et al. (2021); Zhou et al. (2021); Yan et al. (2022b) to obtain both the dimer and vison spectral functions in real frequency for $L=6,12$ systems from imaginary time correlation functions with $\tau\in[0,\beta=200]$.
@@
where $\langle t_{\alpha}(0)t_{\alpha}(\mathbf{r})\rangle$ $(\alpha=1,2,3)$ represent correlators on the three rhombus directions in our triangular lattice with distance $\mathbf{r}$ between two rhombi. The reason of discarding the off-diagonal terms in Eq. (2) will be explained below Eq. (5). The vison correlation function, constructed from the dimer configurations, is
@@
As mentioned above, the $t$-term operator $t_{i}$ can be identified as the field theory operators, $\{t_{1},t_{2},t_{3}\}\sim\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$. The symmetry group of the CFT is the Cubic(3)=$S_{3}\rtimes(Z_{2})^{3}$ group, the group elements of Cubic(3) can be identified with lattice symmetries. The precise identification of $t$-operators is fixed by the symmetries that they break. Here we are following the convention of Ref. Yan et al. (2022a). The Cubic(3) group is a subgroup of O(3). It is known, based on various theoretical works Ballesteros et al. (1996); Calabrese et al. (2003); Chester et al. (2021), that the O(3) CFT and the Cubic(3) are connected by a very short renormalizations group flow, therefore their operators have similar anomalous dimensions. In particular, the O(3) group has a rank-2 symmetric traceless tensor representation, formed by $\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$ and $\{\phi_{1}^{2}-\phi_{2}^{2},\phi_{2}^{2}-\phi_{3}^{2}\}$, which is five dimensional. In the view of the subgroup Cubic(3), the triple $\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$ forms a three dimensional irreducible representation of the Cubic(3) group. We can safely use the well-known value of the critical exponents $\eta_{T}\approx 1.42$ of O(3) CFT to approximate its value at the Cubic CFT Chester et al. (2021). The subscript “$T$” reminds us that it corresponds to the rank-2 tensor of O(3). Interestingly, the off-diagonal correlator $\langle t_{1}(0)t_{2}(\mathbf{r})\rangle$ decay much faster than the diagonal ones $\langle t_{1}(0)t_{1}(\mathbf{r})\rangle$, which is also a CFT prediction and we show these results in Fig.S1 in the SM sup . The anomalous dimension of scalar $\{\phi_{1},\phi_{2},\phi_{3}\}$ for O(3) CFT, i.e. the vison $v_{1,2}$ correlation in Eq. (3), on the other hand, is of very small value $\eta\approx 0.04$ Ballesteros et al. (1996); Chester et al. (2021).
@@
Figure 2: The static $t$-term correlation. (a) $\langle T(0)T(\mathbf{r})\rangle$ as a function of the distance $\mathbf{r}=(x,0)$ with the system size $L=24$, the largest system size achieved. The log-log plot for the absolute values of the data in (a) is shown in (b). We also show the log-log plot of the $t$-term correlators with different system sizes in (c) to demonstrate the finite-size effect of the decay behavior. In (a) and (b), we show the correlators in the VP phase with $V=0.3$, at the transition point with $V_{c}$ (we use $V=0.6$ here), and at the RK point when $V=1$. The dark solid line shown in (b) and (c) is proportional to $1/x^{1+\eta_{T}}$ with the large anomalous dimension $\eta_{T}=1.42$, which corresponds to the rank-2 tensor field of the (2+1)d O(3)/Cubic universality Chester et al. (2021).
@@
Figure 3: The static vison correlation. (a) $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ as a function of the distance $\mathbf{r}=(x,0)$ with the fixed system size $L=24$. (b) The log-log plot of only the odd value of $x$ in (a). Similar with the $t$-term correlators, We show the vison correlators in the VP phase ($V=0.3$), at the transition point $V_{c}$ (use $V=0.61$ here), and at the RK point when $V=1$. Panel (c) shows the outstanding critical decay behavior with different system sizes. Different from the $t$-term correlators in Figs. 2 (b) and (c), the dark solid line shown in (b) and (c) here is proportional to $1/x^{1+\eta}$ with the anomalous dimension $\eta=0.04$ for the (2+1)d O(3)/Cubic scalar order parameter Ballesteros et al. (1996); Chester et al. (2021).
@@
We also compute the dynamic dimer correlation function $D(\mathbf{k},\tau)=\frac{1}{3N}\sum_{\begin{subarray}{c}i,j;\alpha=1,2,3\end{subarray}}^{L^{2}}e^{i\mathbf{k}\cdot\mathbf{r}_{ij}}\left(\langle n_{i,\alpha}(\tau)n_{j,\alpha}(0)\rangle-\langle n_{i,\alpha}\rangle\langle n_{j,\alpha}\rangle\right)$, where $n_{i,\alpha}$ is the dimer number operator on bond $i$ and $\alpha$ stands for the three bond orientations, and the vison dynamic correlation function $\bar{v}(\mathbf{k},\tau)=\frac{1}{2N}\sum_{\begin{subarray}{c}i,j;\gamma=1,2\end{subarray}}^{L^{2}}e^{i\mathbf{k}\cdot\mathbf{r}_{ij}}\left(\langle v_{i,\gamma}(\tau)v_{j,\gamma}(0)\rangle-\langle v_{i,\gamma}\rangle\langle v_{j,\gamma}\rangle\right)$, which averages the correlation functions of visons in A, B sublattices. Since the value of vison in each triangle is ±$\pm 1$, the second term $\langle v_{i,\gamma}\rangle$ in $\bar{v}(\mathbf{k},\tau)$ is expected to be zero, i.e., no background needs to be subtracted.
@@
Figure 4: The dynamic dimer and vison spectra. The spectra in the VP phase $(V=0.3)$, at the Cubic* QCP $(V=0.6)$, and at the RK point $(V=1)$ for $L=12$ system. The $\beta$ used in the simulations is $100$ and we employ QMC+SAC scheme to generate the real frequency data. The inset in (f) shows the high-symmetry path in the Brillouin zone along which the spectra are presented. For dimer spectra in the left column (a), (b) and (e), all the spectra exhibit continua due to the fractionalization. At the $\mathbf{M}$ point, the spectra are gapped, indicating dimer correlator is unable to detect the transition between the VP and QSL phase. While for vison spectra in the right column (d), (e) and (f), the gap at $\mathbf{M}$ point closes at the Cubic* QCP in panel (e), and reopens at the RK point in panel (f). The vison gap at the RK point in (f) is smaller than that of the dimer gap in (c).
@@
Numerical results.— Figs. 2 and 3 show the $\langle T(0)T(\mathbf{r})\rangle$ and $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ across the Cubic* QCP with system size up to $L=24$. The distance is along the $\mathbf{r}=(x,0)$ with $x$ up to $12$ for the periodic boundary condition. The real-space decay behaviors is observed for both correlators in three regions: (i) VP phase with $V=0.3$. (ii) The Cubic* QCP $V_{c}=0.59(2)$. (iii) The RK point $V=1$.
@@
In the VP phase, both $\langle T(0)T(\mathbf{r})\rangle$ and $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ exhibit strong even-odd oscillations and with amplitude decaying with the distance $x$. The oscillations derive from the hidden vison order and eventually vanish as $V$ goes to 1 as shown in Figs. 2 (a) and 3 (a). We note the even-odd oscillation still exists at the transition point due to the finite-size effect. The oscillations of all $V$ are symmetric with respect to $\langle T(0)T(\mathbf{r})\rangle=0$, therefore, we illustrate $|\langle T(0)T(\mathbf{r})\rangle|$ in log-log scale in Fig. 2 (b) and (c). Moreover, due to the gauge choice we set manually to construct the vison configuration, the oscillations of the vison correlation are asymmetrical with respect to $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle=0$ for different values of $V$. Thus we only use the odd value of the distance to fit the data of $|\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle|$ in log-log scale, as shown in Fig. 3 (b) and (c).

### v3

# Cubic* criticality emerging from a quantum loop model on triangular lattice

@@ Abstract
Quantum loop and dimer models are archetypal examples of correlated systems with local constraints. Obtaining generic solutions for these models is difficult due to the lack of controlled methods to solve them in the thermodynamic limit. Nevertheless, these solutions are of immediate relevance to both statistical and quantum field theories, as well as the rapidly growing experiments in Rydberg atom arrays and quantum moiré materials, where the interplay between correlation and local constraints gives rise to a plethora of novel phenomena. In a recent work [X. Ran, Z. Yan, Y.-C. Wang, et al, arXiv:2205.04472 (2022)], it was found through sweeping cluster quantum Monte Carlo (QMC) simulations and field theory analysis that the triangular lattice quantum loop model (QLM) hosts a rich ground state phase diagram with lattice nematic, vison plaquette (VP) crystals, and the $\mathbb{Z}_{2}$ quantum spin liquid (QSL) close to the Rokhsar-Kivelson point. Here, we focus on the continuous quantum critical point separating the VP and QSL phases and demonstrate via both static and dynamic probes in QMC simulations that this transition is of the (2+1)D cubic* universality. In this transition, the fractionalized visons in QSL condense to give rise to the crystalline VP phase, while leaving their trace in the anomalously large anomalous dimension exponent and pronounced continua in the dimer and vison spectra compared with those at the conventional cubic or O(3) quantum critical points.
@@
Introduction.— Recently, the ground state phase diagram of the quantum loop model (QLM) on a triangular lattice Moessner and Sondhi (2001a); Moessner et al. (2001a); Plat et al. (2015); Roychowdhury et al. (2015) has been mapped out using the sweeping cluster quantum Monte Carlo (QMC) algorithm Yan et al. (2022a, 2019, 2021a); Yan (2022); Yan et al. (2021b, 2022b, 2023) (shown in Fig. 1). The physics revealed therein Yan et al. (2022a) is profound, such as the hidden vison plaquette (VP) crystal, which is invisible from dimer correlations and is sandwiched between the lattice nematic (LN) order and the $\mathbb{Z}_{2}$ quantum spin liquid (QSL) close to the Rokhsar-Kivelson (RK) point Rokhsar and Kivelson (1988); Moessner and Sondhi (2001b); Henley (2004). Another interesting aspect is the structure of the phase diagram when connected to finite temperature, which is expected to be richer compared to its square lattice loop or dimer model cousins Alet et al. (2005a); Dabholkar et al. (2022, 2023). However, perhaps the most intriguing aspect related to the quantum criticality of the model is the (2+1)D cubic* transition from the VP phase to the $\mathbb{Z}_{2}$ QSL. At this transition, the VP order parameter, which emerges from the underlying resonance of the dimer pairs, fractionalizes into the vison order parameter of the O(3)/cubic CFT primary field. The condensation of these fractionalized excitations, in turn, leads to a strong enhancement of the scaling dimension of the VP order parameter at the transition in an unconventional manner Wang et al. (2017); Sun et al. (2018); Wang et al. (2018); Isakov et al. (2012); Wang et al. (2021, 2021).
@@
Therefore, our motivation in this Letter is to elucidate the precise nature of the intriguing and unconventional cubic* quantum critical point (QCP) through both static and dynamic probes. We aim to achieve a comprehensive understanding of this QCP by combining field-theoretical interpretation with state-of-the-art QMC simulations. We find that this transition exhibits an anomalously large anomalous dimension when viewed through the correlation functions of the $t$-term (the resonance term in the QLM Hamiltonian, explained below). These correlation functions represent composite objects of the fractionalized vison and correspond to the rank-2 tensor (or tensorial magnetization) of the (2+1)D O(3)/cubic universality Hasenbusch (2023a, b) with a large scaling dimension, approximately $\eta_{T}\approx 1.42$ Ballesteros et al. (1996); Aharony (1973); Hasenbusch and Vicari (2011); Adzhemyan et al. (2019); Aharony et al. (2022); Pelissetto and Vicari (2002); Chester et al. (2021); Rong and Su (2023). On the other hand, if one measures the correlation of the vison operator, the observed anomalous dimension is consistent with the conventionally small values of $\eta\approx 0.04$ for (2+1)D O(3)/cubic universality Ballesteros et al. (1996); Aharony (1973); Hasenbusch and Vicari (2011); Adzhemyan et al. (2019); Aharony et al. (2022); Pelissetto and Vicari (2002); Chester et al. (2021). This sharp contrast clearly reveals the unconventional nature of the cubic* transition that separates the unconventional VP phase, which is hidden from dimer measurements, from the $\mathbb{Z}_{2}$ QSL, where visons are the anyonic particles of the underlying topological order.
@@
Figure 1: Quantum loop model on the triangular lattice and its phase diagram. (a) $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the triangular lattice primitive vectors. The $t$ and $V$ terms are the kinetic and potential terms in the Hamiltonian Eq. (1), respectively. (b) The transition between the LN and VP crystals is first order Yan et al. (2022a), while the transition between VP and the $\mathbb{Z}_{2}$ QSL is continuous and of (2+1)D cubic* universality Yan et al. (2022a). The correlation functions of the VP and vison order parameters around the cubic* QCP ($V_{c}=0.59(2)$) are shown in Figs. 2 and 3. The schematic plot in the VP phase is the real-space vison correlations with the red (grey) color conveying its positive (negative) value in each triangle. The darker color stands for larger absolute values of the correlation functions. The schematic plot in the QSL phase shows two visons connected by a string (the green dashed line), which represents the path $P$ of the vison-vison correlation function $v_{\gamma}(0)v_{\gamma}(r)$$=$$(-1)^{N_{P}}$, with $N_{P}$ the number of dimers cut along $P$. Here we set the vison in the lower triangle $v_{1}(0)=1$ as the reference to fix the gauge.
@@
In addition to these purely theoretical motivations, the QLM that we studied here has been widely treated as the low-energy effective model for many frustrated magnets Moessner and Sondhi (2001a); Moessner et al. (2001a); Roychowdhury et al. (2015); Plat et al. (2015); Yan et al. (2021b, 2022b, 2023); Wang et al. (2017, 2017); Zhang et al. (2018); Feng et al. (2017); Wen (2017); Wei et al. (2020); Li et al. (2020); Hu et al. (2020); Wen et al. (2019); Feng et al. (2018); Wen and Lee (2019); Zhou et al. (2017); Broholm et al. (2020); Da Liao et al. (2021); Zhou et al. (2022); Moessner et al. (2001b); Ivanov (2004); Ralko et al. (2005); Misguich and Mila (2008); Ran et al. (2023) and blockaded cold-atom arrays  Samajdar et al. (2021); Verresen et al. (2021); Zhou et al. ; Samajdar et al. (2023); Semeghini et al. (2021) in condensed matter and cold atom experiments Semeghini et al. (2021); Scholl et al. (2021). In the Rydberg array, static characteristics can be easily obtained via the snapshot technique Scholl et al. (2021); Semeghini et al. (2021); Tran et al. (2023); Wurtz et al. (2023), while dynamic information can be measured through real-time evolution Surace et al. (2020); Notarnicola et al. (2020); Guardado-Sanchez et al. (2018); Turner et al. (2018). Similarly, static and dynamic information for quantum magnets can be detected by neutron scattering or nuclear magnetic resonance experiments Feng et al. (2017); Wei et al. (2020); Hu et al. (2020); Paddison et al. (2017); Zorko et al. (2017); Wen et al. (2019); Zeng et al. (2023), and our computational scheme of QMC + stochastic analytic continuation (SAC) Sandvik (1998a); Beach (2004); Syljuåsen (2008); Shao and Sandvik (2023) for the frustrated spin model, QDM, and QLM models has provided consistent static and dynamic information that has been used to explain experiments Shao et al. (2017); Li et al. (2020); Hu et al. (2020); Sun et al. (2018); Ma et al. (2018); Wang et al. (2021, 2021); Zhou et al. (2021); Yan et al. (2021b, 2022b, 2023); Bera et al. (2022). Based on these previous experiences, in this Letter, our QMC static correlations reveal different scaling dimensions at the cubic* QCP, corresponding to the different constituent operators in the CFT data for $\eta_{T}$ and $\eta$. At the same time, our QMC+SAC dynamic measurements exhibit continua of the dimer and vison spectra as the dynamic signature of the $\mathbb{Z}_{2}$ topological order and its associated vison condensation in the vicinity of the cubic* QCP.
@@
where $\alpha$ denotes all the rhombi (with three orientations) on the triangular lattice, as shown in Fig. 1(a). The local constraint of the fully packed QLM requires two dimers to touch every site in any configuration. The kinetic term is controlled by $t$, which generates dimer pair resonance on every flippable plaquette while respecting the local constraint, and $V$ is the repulsion ($V>0$) or attraction ($V<0$) between dimers facing each other. The RK point is located at $V=t=1$ and has an exact $\mathbb{Z}_{2}$ QSL solution Moessner and Sondhi (2001a). We set $t=1$ as the energy unit and perform simulations for system sizes $L=8,12,16,20,24$ with the inverse temperature $\beta=\frac{1}{T}=L$ using the sweeping cluster QMC methods Yan et al. (2019); Yan (2022); Yan et al. (2021b, 2022b, 2022), and utilize the SAC scheme Sandvik (1998b); Shao et al. (2017); Sun et al. (2018); Ma et al. (2018); Yan et al. (2021b); Shao and Sandvik (2023); Li et al. (2020); Hu et al. (2020); Wang et al. (2021); Zhou et al. (2021); Yan et al. (2022b) to obtain both the dimer and vison spectral functions in real frequency for $L=6,12$ systems from imaginary time correlation functions with $\tau\in[0,\beta=200]$.
@@
where $\langle t_{\alpha}(0)t_{\alpha}(\mathbf{r})\rangle$ $(\alpha=1,2,3)$ represent correlators on the three rhombus directions in our triangular lattice with distance $\mathbf{r}$ between two rhombi. The reason for discarding the off-diagonal terms in Eq. (2) will be explained below Eq. (5). The vison correlation function, constructed from the dimer configurations, is
@@
As mentioned above, the $t$-term operator $t_{i}$ can be identified as the field theory operators, $\{t_{1},t_{2},t_{3}\}\sim\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$. The symmetry group of the CFT is the cubic(3)=$S_{3}\rtimes(Z_{2})^{3}$ group, the group elements of cubic(3) can be identified with lattice symmetries. The precise identification of $t$-operators is fixed by the symmetries that they break. Here we are following the convention of Ref. Yan et al. (2022a). The cubic(3) group is a subgroup of O(3). It is known, based on various theoretical works Ballesteros et al. (1996); Calabrese et al. (2003); Chester et al. (2021), that the O(3) CFT and the cubic(3) are connected by a very short renormalizations group flow, therefore their operators have similar anomalous dimensions. In particular, the O(3) group has a rank-2 symmetric traceless tensor representation, formed by $\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$ and $\{\phi_{1}^{2}-\phi_{2}^{2},\phi_{2}^{2}-\phi_{3}^{2}\}$, which is five-dimensional. In view of the subgroup cubic(3), the triple $\{\phi_{1}\phi_{2},\phi_{2}\phi_{3},-\phi_{1}\phi_{3}\}$ forms a three-dimensional irreducible representation of the cubic(3) group. We can safely use the well-known value of the critical exponents $\eta_{T}\approx 1.42$ of O(3) CFT to approximate its value at the cubic CFT Chester et al. (2021). The subscript “$T$” reminds us that it corresponds to the rank-2 tensor of O(3). Interestingly, the off-diagonal correlator $\langle t_{1}(0)t_{2}(\mathbf{r})\rangle$ decays much faster than the diagonal ones $\langle t_{1}(0)t_{1}(\mathbf{r})\rangle$, which is also a CFT prediction and we show these results in Fig.S1 in the SM sup . The anomalous dimension of scalar $\{\phi_{1},\phi_{2},\phi_{3}\}$ for O(3) CFT, i.e. the vison $v_{1,2}$ correlation in Eq. (3), on the other hand, is of very small value $\eta\approx 0.04$ Ballesteros et al. (1996); Chester et al. (2021).
@@
Figure 2: The static $t$-term correlation. (a) $\langle T(0)T(\mathbf{r})\rangle$ as a function of the distance $\mathbf{r}=(x,0)$ with the system size $L=24$, the largest system size achieved. The log-log plot for the absolute values of the data in (a) is shown in (b). We also show the log-log plot of the $t$-term correlators with different system sizes in (c) to demonstrate the finite-size effect of the decay behavior. (a), (b) The correlators in the VP phase with $V=0.3$, at the transition point with $V_{c}$ (we use $V=0.6$ here), and at the RK point when $V=1$. The dark solid lines shown in (b) and (c) are proportional to $1/x^{1+\eta_{T}}$ with the large anomalous dimension $\eta_{T}=1.42$, which corresponds to the rank-2 tensor field of the (2+1)D O(3)/cubic universality Chester et al. (2021).
@@
Figure 3: The static vison correlation. (a) $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ as a function of the distance $\mathbf{r}=(x,0)$ with the fixed system size $L=24$. (b) The log-log plot of only the odd value of $x$ in (a). Similar to the $t$-term correlators, we show the vison correlators in the VP phase ($V=0.3$), at the transition point $V_{c}$ (use $V=0.61$ here), and at the RK point when $V=1$. (c) The outstanding critical decay behavior with different system sizes. Different from the $t$-term correlators in Figs. 2 (b) and  2(c), the dark solid lines shown in (b) and (c) here are proportional to $1/x^{1+\eta}$ with the anomalous dimension $\eta=0.04$ for the (2+1)D O(3)/cubic scalar order parameter Ballesteros et al. (1996); Chester et al. (2021).
@@
We also compute the dynamic dimer correlation function $D(\mathbf{k},\tau)=\frac{1}{3N}\sum_{\begin{subarray}{c}i,j;\alpha=1,2,3\end{subarray}}^{L^{2}}e^{i\mathbf{k}\cdot\mathbf{r}_{ij}}\left(\langle n_{i,\alpha}(\tau)n_{j,\alpha}(0)\rangle-\langle n_{i,\alpha}\rangle\langle n_{j,\alpha}\rangle\right)$, where $n_{i,\alpha}$ is the dimer number operator on bond $i$ and $\alpha$ stands for the three bond orientations, and the vison dynamic correlation function $\bar{v}(\mathbf{k},\tau)=\frac{1}{2N}\sum_{\begin{subarray}{c}i,j;\gamma=1,2\end{subarray}}^{L^{2}}e^{i\mathbf{k}\cdot\mathbf{r}_{ij}}\left(\langle v_{i,\gamma}(\tau)v_{j,\gamma}(0)\rangle-\langle v_{i,\gamma}\rangle\langle v_{j,\gamma}\rangle\right)$, which averages the correlation functions of visons in A and B sublattices. Since the value of vison in each triangle is ±$\pm 1$, the second term $\langle v_{i,\gamma}\rangle$ in $\bar{v}(\mathbf{k},\tau)$ is expected to be zero, i.e., no background needs to be subtracted.
@@
Figure 4: The dynamic dimer and vison spectra. The spectra in the VP phase $(V=0.3)$, at the cubic* QCP $(V=0.6)$, and at the RK point $(V=1)$ for $L=12$ system. The $\beta$ used in the simulations is $100$ and we employ QMC+SAC scheme to generate the real frequency data. The inset in (f) shows the high-symmetry path in the Brillouin zone along which the spectra are presented. In the dimer spectra displayed in the left column, panels (a), (b), and (c), a gap is observed at the $\mathbf{M}$ point, suggesting that the dimer correlator cannot detect the transition between the VP and QSL phases. Conversely, the vison spectra in the right column, panels (d), (e), and (f), reveal a gap closure at the $\mathbf{M}$ point at the cubic* QCP in panel (e) and a reopening at the RK point in (f), which clearly indicates the VP-QSL transition.
@@
Numerical results.— Figures. 2 and 3 show the $\langle T(0)T(\mathbf{r})\rangle$ and $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ across the cubic* QCP with system size up to $L=24$. The distance is along $\mathbf{r}=(x,0)$ with $x$ up to $12$ for the periodic boundary condition. The real-space decay behaviors is observed for both correlators in three regions: (i) VP phase with $V=0.3$. (ii) The cubic* QCP $V_{c}=0.59(2)$. (iii) The RK point $V=1$.
@@
In the VP phase, both $\langle T(0)T(\mathbf{r})\rangle$ and $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle$ exhibit strong even-odd oscillations and with amplitude decaying with the distance $x$. The oscillations derive from the hidden vison order and eventually vanish as $V$ goes to 1 as shown in Figs. 2 (a) and 3 (a). We note the even-odd oscillation still exists at the transition point due to the finite-size effect. The oscillations of all $V$ are symmetric with respect to $\langle T(0)T(\mathbf{r})\rangle=0$, therefore, we illustrate $|\langle T(0)T(\mathbf{r})\rangle|$ in log-log scale in Figs. 2 (b) and  2(c). Moreover, due to the gauge choice we set manually to construct the vison configuration, the oscillations of the vison correlation are asymmetrical with respect to $\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle=0$ for different values of $V$. Thus, we only use the odd value of the distance to fit the data of $|\langle\bar{v}(0)\bar{v}(\mathbf{r})\rangle|$ in log-log scale, as shown in Figs. 3 (b) and  3(c).